When an object is moving along the principal axis of a lens at object distances 8 cm and 16 cm , the images being thrice the size of the object can be found. Find the type and focal length of the lens
1
Expert's answer
2020-10-29T07:02:36-0400

We have object distance of -16 cm and the magnification is -3. Since magnification is


"m=\frac{i}{o}\rightarrow i=mo=(-16)(-3)=48\text{ cm}."

According to thin lens equation:


"\frac{1}f=\frac{1}i+\frac{1}o,\space\space\space\space\space\space\frac{1}f=\frac{1}{48}+\frac{1}{(-16)},\\\space\\
f=12\text{ cm}."

This is a convex lens because a concave (diverging) lens cannot produce magnified images (i.e. magnification greater than unity).